OpenStudy (anonymous):

Mrs. Hernanadez is covering the top of a kitchen shelf that is 26 inches long and 15 inches wide. She incorrectly measures the length to be 25 inches long. How will this error affect her calculations? a. It will cause her to overestimate the shelf by about 25 square inches c. It will cause her to overestimate the shelf by about 21 square inches b. It will cause her to underestimate the shelf by about 25 square inches d. It will cause her to underestimate the shelf by about 21 square inches

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

If the real area of the top of the kitchen shelf is A=26*15 square inches=390 square inches, what you have to do is calculate the incorrect area which is A=25*15 square inches= 375 square inches, then subtract the incorrect area from the correct one. 390 square inches - 375 square inches = 15 square inches. This means that she will underestimate the shelf by about 15 square inches. This is what comes to my mind. I know that this is not on your answers, but I hope you find this helpful.